OpenStudy (anonymous):

a=p(1+r) a=82.5 AND P=75 THEN R=

OpenStudy (anonymous):

82.5 = 75(1+r) 82.5/75 = 1+r 82.5/75 -1 = r 0.1 = r

jhonyy9 (jhonyy9):

7,5/75 =r